OpenStudy (anonymous):

Write the equation of the line given the slope and the y-intercept, m = -4, point (2, -5)

OpenStudy (anonymous):

(2,-5) is not a intercept

OpenStudy (anonymous):

y=mx+c -5=(-4)(2)+c 8-5=c 3=c y=-4c+3 ur answer